A) Calculate the number of molecules of CH4 in 72 g CH4.

B) The number of grams in 3.33 mol of sodium carbonate is

C) 3.44 g of Ca represents how many moles?

D) 8.55 x 10-3 moles of an unknown compound has a mass of 0.376 g. The compound could be:
Answer #1

a.

number of moles of CH4 = 72 g / 16 g/mol = 4.5 mol

number of molecules = 4.5 mol * (6.022 * 1023 molecules / mole) = 2.71 * 1024 molecules

b.

mass = number of moles * molar mass = 3.33 mol * 106 g/mol = 352.98 g

c.

number of moles = 3.44 g / 40 g/mol = 0.086 moles

d.

molar mass = mass / number of moles = (0.376g / 8.55*10-3 mol) = 43.97 g/mol